OpenStudy (anonymous):

What is the equation of a line that passes through (7, 8) and has a slope of -3? a. y=-3x+29 b. y=3x+13 c. y=1/3x-29 d. y=-1/3x-13

OpenStudy (freckles):

y=mx+b is slope-intercept form m is the slope and b is the y-intercept

OpenStudy (freckles):

we are given m we are also given a point (x,y) on the line enter your point given and your value for m given into y=mx+b and solve for b

OpenStudy (anonymous):

b?

OpenStudy (freckles):

remember y=mx+b is in slope-intercept form it tells you the slope is m and the y-intercept is b so compare y=3x+13 to y=mx+b what is m here?

OpenStudy (anonymous):

3